OpenStudy (bloomlocke367):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (-5,0) and co-vertex at (0,4)

OpenStudy (bloomlocke367):

I'll fan and metal!!!!!!!!!!!

OpenStudy (bloomlocke367):

@SolomonZelman

OpenStudy (bloomlocke367):

is it: \[\frac{ x^2 }{ 25 }+\frac{ y^2 }{ 16 }=1\]

OpenStudy (bloomlocke367):

@phi @ganeshie8 @hartnn @iambatman

ganeshie8 (ganeshie8):

Yep!

OpenStudy (bloomlocke367):

ok.. can you help me more?

ganeshie8 (ganeshie8):

tag me.. il try :) also wolfram is a good friend for double checkign any kind of conics problems : http://www.wolframalpha.com/input/?i=vertex+x%5E2%2F25+%2B+y%5E2%2F16%3D1

ganeshie8 (ganeshie8):

scroll all the way down to see other properties of that ellipse..

OpenStudy (bloomlocke367):

ok... thanks